# The KSBA moduli space of stable log Calabi–Yau surfaces

### v2

@@ 4. Semistable mirrors of polarized log Calabi–Yau surfaces
In this section, we construct in §4.2 a quasi-projective open Kulikov surface $\mathcal{X}_{\mathscr{P},0}$, associated to a polarized log Calabi–Yau $(Y,D,L)$ and a choice of Symington polytope $P$ endowed with a good polyhedral decomposition $\mathscr{P}$. We describe $\mathcal{X}_{\mathscr{P},0}$ as a union of suitable “deformations” of toric varieties determined by the polyhedral decomposition $\mathscr{P}$. The necessary results in deformation theory of log Calabi–Yau surfaces are first established in §4.1. Finally, we define in §4.3 the semistable mirror to $(Y,D,L)$ as an open Kulikov degeneration $\mathcal{X}_{\mathscr{P}}\rightarrow\Delta$ obtained as the total space of a semistable smoothing of the open Kulikov surface $\mathcal{X}_{\mathscr{P},0}$.
@@ 4.1. Deformation theory of log Calabi–Yau surfaces
The result is well-known if $D$ is empty, see for example [82, Theorem 4.10], [76, Lemma 1], or [97, Proposition 6.4]. On the other hand, if $D$ is non-empty but $X$ is smooth, a proof can be found in [34, Proposition 3.5]. Our proof below of the general case is based on a combination of the techniques used for these two known special cases.
Let $U$ be the non-empty open subset of $X$ where both $X$ and $D$ are smooth, and denote by $j:U\hookrightarrow X$ the inclusion. By assumption, the complement $X\setminus D$ consists of isolated points in the surface $X$, and so if of codimension two. Consider the complex $A:=(j_{\star}T_{X}|_{U}\rightarrow j_{\star}N_{D|U})$, where $T_{X}|_{U}$ is the restriction to $U$ of the tangent sheaf $T_{X}$, placed in degree 00, and $N_{D|U}=\mathcal{O}_{D\cap U}(D\cap U)$ is the normal sheaf to $D\cap U$ in $D\cap U$, placed in degree $1$. Using as in the proof of [82, Theorem 4.10] a Čech cover of $X$ by affine subsets centered on the isolated singularities and with double and triple intersections contained in $U$, we deduce that if the hypercohomology group $\mathbb{H}^{2}(X,A)$ vanishes, then there are no local-to-global obstructions to deformations of $(X,D)$. When $X$ is smooth, details of this argument can be found in [88, Proposition 8].
@@
defining $D\cap U$ in $U$. As $T_{X}|_{U}$ is locally free, tensoring by $T_{X}|_{U}$, and then applying $j_{\star}$, we obtain an exact sequence of the form
@@
where $P$ is a sheaf supported on $X\setminus U$. Note that as $T_{X}|_{U}(-D\cap U)$ is locally free on $U$, we have $j_{\star}T_{X}|_{U}(-D\cap U)=(T_{X}(-D))^{\vee\vee}$. We split the above long exact sequence into two short exact sequences:
@@
Hence, it is enough to show that $\mathbb{H}^{2}(X,B)=0$ in order to show the desired result that $H^{2}(X,(T_{X}(-D))^{\vee\vee})=0$ implies $\mathbb{H}^{2}(X,A)=0$.
@@
The following argument is inspired by the proof of [76, Theorem 21]. By Lemma 4.1, it is enough to show that $H^{2}(X,(T_{X}(K_{X}))^{\vee\vee})=0$. A normal surface being Cohen-Macaulay, the reflexive canonical sheaf $\mathcal{O}_{X}(K_{X})$ is the dualizing sheaf of $X$, and so by Serre duality we have
@@
Let $j:U\hookrightarrow X$ be the open immersion of the smooth locus of $X$. As $X$ is normal and $X\setminus U$ is of codimension two, the functors $j_{\star}$ and $j^{\star}$ are inverse equivalences between the category of coherent reflexive sheaves on $X$ and the coherent reflexive sheaves on $U$ [89, Lemma 0EBJ]. As both $(T_{X}(K_{X}))^{\vee\vee}$ and $\mathcal{O}_{X}(K_{X})$ are reflexive, it follows that
@@
Finally, as $X$ has quotient singularities, it follows from Steenbrink Hodge theory for orbifolds [90, §1] that $\dim H^{0}(X,\Omega_{X}^{\vee\vee})=\dim H^{1}(X,\mathcal{O}_{X})$. Hence, it follows that
@@
4.2. Open Kulikov surfaces from polarized log Calabi-Yau surfaces
Let $(Y,D,L)$ be a polarized log Calabi–Yau surface and $P$ a Symington polytope of $(Y,D,L)$ defined by a good toric model $(\overline{Y},\overline{D},\overline{L})$ with toric momentum polytope $\overline{P}$, as in §2. In this section, we describe the construction, for every good polyhedral decomposition $\overline{\mathscr{P}}$ of $\overline{P}$ as in §2.3, of a quasi-projective open Kulikov surface $\mathcal{X}_{\mathscr{P},0}$, which is unique up to locally trivial deformations.
Throughout this section we denote by $M\cong\mathbb{Z}^{2}$ the character lattice for $\overline{Y}$, so that $\overline{P}\subseteq M_{\mathbb{R}}$, where $M_{\mathbb{R}}=M\otimes_{\mathbb{Z}}\mathbb{R}$ denotes the real vector space associated to $M$. We first consider the Mumford toric degeneration $\pi_{\overline{\mathscr{P}}}:\mathcal{X}_{\overline{\mathscr{P}}}\rightarrow\mathbb{A}^{1}$ associated to the toric polytope $\overline{P}$ endowed with the polyhedral decomposition $\overline{\mathscr{P}}$. As in [78, §3], $\mathcal{X}_{\overline{\mathscr{P}}}$ is the 3-dimensional toric variety with fan $C(\overline{\mathscr{P}})$ in $M_{\mathbb{R}}\times\mathbb{R}_{\geq 0}$ consisting of the cones
@@ 4.2. Open Kulikov surfaces from polarized log Calabi-Yau surfaces
Since all cells $\sigma\in\overline{\mathscr{P}}$ are bounded, the so called asymptotic cone in [78] consists of a single point, the origin in $M_{\mathbb{R}}\times\mathbb{R}$, and hence the fibers $\pi_{\overline{\mathscr{P}}}^{-1}(t)$ for $t\neq 0$ are isomorphic to the algebraic torus (*$(\mathbb{C}^{*})^{2}$. The central fiber $\mathcal{X}_{\overline{\mathscr{P}},0}\coloneqq\pi_{\overline{\mathscr{P}}}^{-1}(0)$ is a union of irreducible components $X_{\overline{\mathscr{P}}}^{v}$, indexed by the vertices $v\in\overline{\mathscr{P}}^{[0]}$, which are toric surfaces with fan the star of $v$ in $\overline{\mathscr{P}}$:
@@
We consider below the Symington polytope $P$, endowed with the polyhedral decomposition $\mathscr{P}$ induced by the restriction of $\overline{\mathscr{P}}$ to $P$, and we describe a union $\mathcal{X}_{\mathscr{P},0}$ of log Calabi–Yau surfaces obtained as a deformation of $\mathcal{X}_{\overline{\mathscr{P}},0}$, and with irreducible components $\mathcal{X}_{\mathscr{P},0}^{v}$ indexed by the vertices $v\in\mathscr{P}^{[0]}$. We do this in four steps.
@@
STEP I: As a first step, for every vertex $v\in\mathscr{P}^{[0]}$, we define a toric surface $\widetilde{X}^{v}_{\mathscr{P}}$ by deformation of the union $\bigcup_{v^{\prime}\in pr^{-1}(v)}X^{v^{\prime}}_{\overline{\mathscr{P}}}$ of irreducible components of $\mathcal{X}_{\overline{\mathscr{P}},0}$, where $pr:\overline{\mathscr{P}}^{[0]}\rightarrow\mathscr{P}^{[0]}$ is the natural projection coming from the description of the Symington of $P$ as a quotient of $\overline{P}\setminus\cup_{ij}\mathrm{Int}(\Delta_{ij})$. If $pr^{-1}(v)$ is of cardinality one, with $pr^{-1}(v)=\{v^{\prime}\}$, then we set simply $\widetilde{X}^{v}_{\mathscr{P}}:=X^{v^{\prime}}_{\overline{\mathscr{P}}}$. If $pr^{-1}(v)$ is of cardinality strictly bigger than one, then $pr^{-1}(v)=\{v^{\prime}_{1},\dots,v^{\prime}_{r}\}$, where for every $11\leq k\leq r-$ 1, $v^{\prime}_{k}$ and $v^{\prime}_{k+1}$ are opposite integral points on the interior edges of a triangle $\Delta_{i_{k}j_{k}}$. As the good decomposition $\overline{\mathscr{P}}$ is a standard decomposition in restriction to $\Delta_{ij}$, as in Definition 2.13, it follows that each toric surface $X^{v_{k}^{\prime}}_{\overline{\mathscr{P}}}$ has a map to $\mathbb{P}^{1}$ induced by the opposite rays in its fan coming from the interior edge of $\Delta_{i_{k}j_{k}}$ or $\Delta_{i_{k-1}j_{k-1}}$ containing $v_{k}^{\prime}$. These maps $f_{k}:X^{v_{k}^{\prime}}_{\overline{\mathscr{P}}}\rightarrow\mathbb{P}^{1}$ are generically $\mathbb{P}^{1}$-fibrations if $v_{k}^{\prime}$ is contained in the interior of $\overline{P}$, and generically $\mathbb{A}^{1}$-fibrations if $v_{k}^{\prime}$ is on the boundary of $\overline{P}$. Moreover, for every $11\leq k\leq r-$ 1, there is an edge in $\mathscr{P}^{[1]}$ contained in $\Delta_{i_{k}j_{k}}$ and connecting the vertices $v_{k}^{\prime}$ and $v_{k+1}^{\prime}$. This edge corresponds to the intersection $F_{k}:=X^{v_{k}^{\prime}}_{\overline{\mathscr{P}}}\cap X^{v_{k+1}^{\prime}}_{\overline{\mathscr{P}}}$, and is both a fiber of $f_{k}$ in $X^{v_{k}^{\prime}}_{\overline{\mathscr{P}}}$ and a fiber of $f_{k+1}$ in $X^{v_{k+1}^{\prime}}_{\overline{\mathscr{P}}}$.
[…]
@@
Figure 4.1. Deformation of the toric intersection complex of $\mathcal{X}_{\overline{\mathscr{P}},0}$ into the non-toric intersection complex of $\mathcal{X}_{\mathscr{P},0}$.
Hence, $\bigcup_{i=1}^{k}X^{v_{k}^{\prime}}_{\overline{\mathscr{P}}}$ is a chain of fibered surfaces, glued along common fibers. The double locus of such a normal crossing union of toric surfaces can be smoothed by a toric degeneration defined by a polyhedral decomposition of $\mathbb{R}^{2}$ as in Figure 4.1. We define $\widetilde{\mathcal{X}}_{\mathscr{P}}^{v}$ as the toric surface given by the general fiber of this deformation. The fan of $\widetilde{\mathcal{X}}_{\mathscr{P}}^{v}$ can be described explicitly as follows. First we draw the fan of $X^{v_{1}^{\prime}}_{\overline{\mathscr{P}}}$ such that the opposite rays defining $f_{1}$ are horizontal and such that the ray corresponding to $F_{1}$ is contained in the upper-half plane. We also draw the fan of $X^{v_{r}^{\prime}}_{\overline{\mathscr{P}}}$ such that the opposite rays defining $f_{r}$ are horizontal and such that the ray corresponding to $F_{r-1}$ is contained in the lower-half plane and parallel to the ray corresponding to $F_{1}$ in the fan of $X^{v_{1}^{\prime}}_{\overline{\mathscr{P}}}$. Then, the fan of $\widetilde{\mathcal{X}}_{\mathscr{P}}^{v}$ is equal to the fan of $X^{v_{1}^{\prime}}_{\overline{\mathscr{P}}}$ in restriction to the lower half-plane, and equal to the fan of $X^{v_{r}^{\prime}}_{\overline{\mathscr{P}}}$ in restriction to the upper half-plane.
@@
STEP II: As a second step, for each vertex $v\in\mathscr{P}^{[0]}$, we describe a deformation of the toric pair $(\widetilde{X}_{\mathscr{P}}^{v},\partial\widetilde{X}_{\mathscr{P}}^{v})$ into a log Calabi–Yau pair $(X_{\mathscr{P}}^{v},\partial X_{\mathscr{P}}^{v})$. Recall by Construction 2.10, the Symington polytope $P$ is obtained from $\overline{P}$, by cutting out triangles $\Delta_{ij}$, with one edge in the boundary of $\overline{P}$ and with an interior vertex $p_{ij}$. For every $i,j$ such that $v=pr(p_{ij})$, there is a face $F_{ij}$ of $\overline{\mathscr{P}}$ containing $v$ and contained in the triangle $\Delta_{ij}$. The face $F_{ij}$ defines a 2-dimensional cone in the star of $\overline{\mathscr{P}}$ at $v$, that is, in the fan of $X_{\overline{\mathscr{P}}}^{v}$, and so corresponds to a toric fixed point $x_{ij}$ of $X_{\overline{\mathscr{P}}}^{v}$. This point is not smoothed by the deformation in Step I, and so also defines a toric fixed point in the toric surface $\widetilde{X}_{\mathscr{P}}^{v}$ that we still denote by $x_{ij}$. We denote by $(V_{ij},\partial V_{ij})$ the analytic germ of the surface pair $(\widetilde{X}_{\mathscr{P}}^{v},\partial\widetilde{X}_{\mathscr{P}}^{v})$ around $x_{ij}$. As the triangles $\Delta_{ij}$ have equal lattice height and lattice base, it follows from [11, §7.3] that $(V_{ij},\partial V_{ij})$ is the germ of the origin in an affine toric variety of the form
@@
[…]
@@
Taking $\{z=0\}$ as deformation of $\partial V_{ij}$, we view this smoothing of $V_{ij}$ as a one-parameter deformation of the pair $(V_{ij},\partial{V}_{ij})$. If $n=1$, we consider the one-parameter deformation of $(\mathbb{A}^{2}_{u,v},\{uv=0\})$ given by $(\mathbb{A}^{2}_{u,v},\{uv=t\})$, that is by just smoothing the normal crossing toric divisor. Hence, for every value of $n$, we have a natural one-parameter $\mathbb{Q}$-Gorenstein deformation of $(V_{ij},\partial V_{ij})$, whose general fiber is the germ of a smooth point with a smooth connected divisor.

### v3

@@ 4. Semistable mirrors of polarized log Calabi–Yau surfaces
In this section, we construct in §4.2 a quasi-projective open Kulikov surface $\mathcal{X}_{\mathscr{P},0}$, associated to a generic polarized log Calabi–Yau $(Y,D,L)$ and a choice of Symington polytope $P$ endowed with a good polyhedral decomposition $\mathscr{P}$. We describe $\mathcal{X}_{\mathscr{P},0}$ as a union of suitable “deformations” of toric varieties determined by the polyhedral decomposition $\mathscr{P}$. The necessary results in deformation theory of log Calabi–Yau surfaces are first established in §4.1. Finally, we define in §4.3 the semistable mirror to $(Y,D,L)$ as an open Kulikov degeneration $\mathcal{X}_{\mathscr{P}}\rightarrow\Delta$ obtained as the total space of a semistable smoothing of the open Kulikov surface $\mathcal{X}_{\mathscr{P},0}$.
@@ 4.1. Deformation theory of log Calabi–Yau surfaces
The result is well-known if $D$ is empty, see for example [87, Theorem 4.10], [81, Lemma 1], or [101, Proposition 6.4]. On the other hand, if $D$ is non-empty but $X$ is smooth, a proof can be found in [36, Proposition 3.5]. Our proof below of the general case is based on a combination of the techniques used for these two known special cases.
Let $U$ be the non-empty open subset of $X$ where both $X$ and $D$ are smooth, and denote by $j:U\hookrightarrow X$ the inclusion. By assumption, the complement $X\setminus U$ consists of isolated points in the surface $X$, and so is of codimension two. Consider the complex $A:=(j_{\star}T_{X}|_{U}\rightarrow j_{\star}N_{D|U})$, where $T_{X}|_{U}$ is the restriction to $U$ of the tangent sheaf $T_{X}$, placed in degree 0, and $N_{D|U}=\mathcal{O}_{D\cap U}(D\cap U)$ is the normal sheaf to $D\cap U$ in $U$, placed in degree $1$. Using as in the proof of [87, Theorem 4.10] a Čech cover of $X$ by affine subsets centered on the isolated singularities and with double and triple intersections contained in $U$, we deduce that if the hypercohomology group $\mathbb{H}^{2}(X,A)$ vanishes, then there are no local-to-global obstructions to deformations of $(X,D)$. When $X$ is smooth, details of this argument can be found in [92, Proposition 8].
@@
defining $D\cap U$ in $U$. Since $T_{X}|_{U}$ is locally free, we obtain by tensoring with $T_{X}|_{U}$ and then applying $j_{\star}$, an exact sequence of the form
@@
where $P$ is a sheaf supported on $X\setminus U$. Since $T_{X}|_{U}(-D\cap U)$ is locally free on $U$, we have $j_{\star}T_{X}|_{U}(-D\cap U)=(T_{X}(-D))^{\vee\vee}$. We split the above long exact sequence into two short exact sequences:
@@
Hence, it suffices to show that $\mathbb{H}^{2}(X,B)=0$ in order to show the desired result that $H^{2}(X,(T_{X}(-D))^{\vee\vee})=0$ implies $\mathbb{H}^{2}(X,A)=0$.
@@
The following argument is inspired by the proof of [81, Theorem 21]. By Lemma 4.1, it suffices to show that $H^{2}(X,(T_{X}(K_{X}))^{\vee\vee})=0$. A normal surface being Cohen-Macaulay, the reflexive canonical sheaf $\mathcal{O}_{X}(K_{X})$ is the dualizing sheaf of $X$, and so by Serre duality we have
@@
Let $j:U\hookrightarrow X$ be the open immersion of the smooth locus of $X$. Since $X$ is normal and $X\setminus U$ is of codimension two, the functors $j_{\star}$ and $j^{\star}$ are inverse equivalences between the category of coherent reflexive sheaves on $X$ and the coherent reflexive sheaves on $U$ [93, Lemma 0EBJ]. As both $(T_{X}(K_{X}))^{\vee\vee}$ and $\mathcal{O}_{X}(K_{X})$ are reflexive, it follows that
@@
Finally, as $X$ has quotient singularities, it follows from Hodge theory for orbifolds [94, §1] that $\dim H^{0}(X,\Omega_{X}^{\vee\vee})=\dim H^{1}(X,\mathcal{O}_{X})$. Hence, we have
@@
4.2. Open Kulikov surfaces from polarized log Calabi–Yau surfaces
Let $(Y,D,L)$ be a generic polarized log Calabi–Yau surface and $P$ a Symington polytope of $(Y,D,L)$ defined by a good toric model $(\overline{Y},\overline{D},\overline{L})$ with toric momentum polytope $\overline{P}$, as in §2. In this section, we describe the construction, for every good polyhedral decomposition $\overline{\mathscr{P}}$ of $\overline{P}$ as in §2.3, of a quasi-projective open Kulikov surface $\mathcal{X}_{\mathscr{P},0}$, which is unique up to locally trivial deformations.
Throughout this section we denote by $M\cong\mathbb{Z}^{2}$ the character lattice for $\overline{Y}$, so that $\overline{P}\subseteq M_{\mathbb{R}}$, where $M_{\mathbb{R}}=M\otimes_{\mathbb{Z}}\mathbb{R}$ denotes the real vector space associated to $M$. We first consider the Mumford toric degeneration $\pi_{\overline{\mathscr{P}}}:\mathcal{X}_{\overline{\mathscr{P}}}\rightarrow\mathbb{A}^{1}$ associated to the toric polytope $\overline{P}$ endowed with the polyhedral decomposition $\overline{\mathscr{P}}$. As in [83, §3], $\mathcal{X}_{\overline{\mathscr{P}}}$ is the 3-dimensional toric variety with fan $C(\overline{\mathscr{P}})$ in $M_{\mathbb{R}}\times\mathbb{R}_{\geq 0}$ consisting of the cones
@@ 4.2. Open Kulikov surfaces from polarized log Calabi–Yau surfaces
Since all cells $\sigma\in\overline{\mathscr{P}}$ are bounded, the so called asymptotic cone in [83] consists of a single point, the origin in $M_{\mathbb{R}}\times\mathbb{R}$, and hence the fibers $\pi_{\overline{\mathscr{P}}}^{-1}(t)$ for $t\neq 0$ are isomorphic to the algebraic torus $(\mathbb{C}^{*})^{2}$. The central fiber $\mathcal{X}_{\overline{\mathscr{P}},0}\coloneqq\pi_{\overline{\mathscr{P}}}^{-1}(0)$ is a union of irreducible components $X_{\overline{\mathscr{P}}}^{v}$, indexed by the vertices $v\in\overline{\mathscr{P}}^{[0]}$, which are toric surfaces with fan the star of $v$ in $\overline{\mathscr{P}}$:
@@
In this section, we construct by deformation of $\mathcal{X}_{\overline{\mathscr{P}},0}$ a normal crossing surface $\mathcal{X}_{\mathscr{P},0}$, whose irreducible components are log Calabi–Yau surfaces $X_{\mathscr{P}}^{v}$ indexed by the vertices $v\in\mathscr{P}^{[0]}$. The construction proceeds in four steps.
@@
STEP I: Let $pr:\overline{\mathscr{P}}^{[0]}\rightarrow\mathscr{P}^{[0]}$ be the natural projection coming from the description of the Symington of $P$ as a quotient of $\overline{P}\setminus\cup_{ij}\mathrm{Int}(\Delta_{ij})$. As a first step, we construct for every vertex $v\in\mathscr{P}^{[0]}$, a toric surface $\widetilde{X}^{v}_{\mathscr{P}}$ by deformation of the union $\bigcup_{v^{\prime}\in pr^{-1}(v)}X^{v^{\prime}}_{\overline{\mathscr{P}}}$ of irreducible components of $\mathcal{X}_{\overline{\mathscr{P}},0}$. When $pr^{-1}(v)$ is of cardinality one, with $pr^{-1}(v)=\{v^{\prime}\}$, we simply set $\widetilde{X}^{v}_{\mathscr{P}}:=X^{v^{\prime}}_{\overline{\mathscr{P}}}$.
When $pr^{-1}(v)$ is of cardinality strictly bigger than one, then $pr^{-1}(v)=\{v^{\prime}_{1},\dots,v^{\prime}_{r}\}$, where for every $1\leq k\leq r-1$, $v^{\prime}_{k}$ and $v^{\prime}_{k+1}$ are opposite integral points on the interior edges of a triangle $\Delta_{i_{k}j_{k}}$. For each $1\leq k\leq r$, let $\Sigma_{k}$ be the fan of $X^{v_{k}^{\prime}}_{\overline{\mathscr{P}}}$, that is, the star of $\overline{\mathscr{P}}$ at $v_{k}^{\prime}$. Let $f_{k}$ be the primitive integral direction vector pointing towards $p_{i_{k}j_{k}}$ of the interior edge of $\Delta_{i_{k}j_{k}}$ containing $v_{k}^{\prime}$. Denote by $e_{k}$ the primitive direction vector of the edge of $\overline{\mathscr{P}}$ connecting $v_{k}^{\prime}$ to $v_{k+1}^{\prime}$, pointing towards $v_{k+1}^{\prime}$. Each fan $\Sigma_{k}$ contains rays of directions $f_{k}$, $e_{k}$ if $k<r$, and $e_{k-1}$ if $k>1$.
Now, let $T_{k}$ be the unique element of $SL(2,\mathbb{Z})$ such that $T_{k}(f_{k+1})=f_{k}$ and $T_{k}(e_{k})=e_{k}$. Define $\widetilde{\Sigma}_{k}$ as the fan obtained by applying $T_{1}\cdots T_{k-1}$ to $\Sigma_{k}$. Then, the ray of direction $f_{k}$ in $\Sigma_{k}$ becomes a ray of direction $f_{1}$ in $\widetilde{\Sigma}_{k}$. Since all these rays are parallel, there exists a polyhedral decomposition $\mathfrak{F}$ of $\mathbb{R}^{2}$, with vertices $w_{1},\dots,w_{r}$, such that the star of $\mathfrak{F}$ at each vertex $w_{k}$ is precisely $\widetilde{\Sigma}_{k}$, and the vertices $w_{k}$ and $w_{k+1}$ are connected by an edge of direction $T_{1}\cdots T_{k-1}(e_{k})$. An explicit example of this construction is given in Example 4.3 below.
The polyhedral decomposition $\mathfrak{F}$ defines a toric degeneration over $\mathbb{A}^{1}$, as in [83, §3]. Since each fan $\widetilde{\Sigma}_{k}$ is obtained from $\Sigma_{k}$ by applying an element of $SL(2,\mathbb{Z})$, it defines the same toric variety $X^{v_{k}^{\prime}}_{\overline{\mathscr{P}}}$. Thus, the central fiber of the toric degeneration is $\bigcup_{k=1}^{r}X^{v_{k}^{\prime}}_{\overline{\mathscr{P}}}$. We define the toric surface $\widetilde{X}^{v}_{\mathscr{P}}$ as the general fiber of the toric degeneration, whose fan is the asymptotic fan of the polyhedral decomposition $\mathfrak{F}$. The toric degeneration effectively smooths out the double locus of $\bigcup_{k=1}^{r}X^{v_{k}^{\prime}}_{\overline{\mathscr{P}}}$, see Figure 6.
@@
Figure 6. Deformation of the toric intersection complex of $\mathcal{X}_{\overline{\mathscr{P}},0}$ into the non-toric intersection complex of $\mathcal{X}_{\mathscr{P},0}$. The black line turning into the dashed-line represents the part of the double locus of $\mathcal{X}_{\overline{\mathscr{P}},0}$ becoming smooth in $\mathcal{X}_{\mathscr{P},0}$.
Example 4.3.
Consider Figure 7, where the three vertices $v^{\prime}_{1}$, $v^{\prime}_{2}$, $v^{\prime}_{3}$ of $\overline{\mathscr{P}}$ are identified to a single vertex $v$ of $\mathscr{P}$. In this example, we have $f_{1}=(0,1)$, $f_{2}=(-1,1)$, $f_{3}=(-1,0)$, and $e_{1}=(1,0)$, $e_{2}=(0,1)$. The matrices $T_{1},T_{2}\in SL(2,\mathbb{Z})$ such that $T_{1}(f_{2})=f_{1}$, $T_{1}(e_{1})=e_{1}$ and $T_{2}(f_{3})=f_{2}$, $T_{2}(e_{2})=e_{2}$ are given by
$T_{1}=\begin{pmatrix}1&1\\ 0&1\end{pmatrix}\,\,\,\,\text{and}\,\,\,\,T_{2}=\begin{pmatrix}1&0\\ -1&1\end{pmatrix}\,$.
The resulting polyhedral decomposition $\mathfrak{F}$ of $\mathbb{R}^{2}$ is represented on the left of Figure 8. The local fan around $w_{1}$ coincides with the local fan around $v_{1}^{\prime}$, whereas the local fan around $w_{2}$ is obtained from the local fan around $v_{2}^{\prime}$ by applying $T_{1}$, and the local fan around $w_{3}$ is obtained from the local fan around $v_{3}^{\prime}$ by applying $T_{1}T_{2}$. The corresponding asymptotic fan, which gives the fan of the toric surface $\widetilde{X}^{v}_{\mathscr{P}}$, is represented on the right of Figure 8.
Figure 7. The polygon $\overline{P}$ and the polyhedral decomposition $\overline{\mathscr{P}}$ in Example 4.3
Figure 8. On the left, the polyhedral decomposition $\mathfrak{F}$ in Example 4.3. On the right, the corresponding asymptotic fan of $\widetilde{X}^{v}_{\mathscr{P}}$.
@@
STEP II: As a second step, for each vertex $v\in\mathscr{P}^{[0]}$, we describe a deformation of the toric pair $(\widetilde{X}_{\mathscr{P}}^{v},\partial\widetilde{X}_{\mathscr{P}}^{v})$ into a log Calabi–Yau pair $(X_{\mathscr{P}}^{v},\partial X_{\mathscr{P}}^{v})$. Recall from Construction 2.10 that the Symington polytope $P$ is obtained from $\overline{P}$ by cutting out triangles $\Delta_{ij}$, each having one edge on the boundary of $\overline{P}$ and an interior vertex $p_{ij}$. For every $i,j$ such that $v=pr(p_{ij})$, there is a face $F_{ij}$ of $\overline{\mathscr{P}}$ containing $v$ and contained in the triangle $\Delta_{ij}$. The face $F_{ij}$ defines a 2-dimensional cone in the star of $\overline{\mathscr{P}}$ at $v$, that is, in the fan of $X_{\overline{\mathscr{P}}}^{v}$, and so corresponds to a toric fixed point $x_{ij}$ of $X_{\overline{\mathscr{P}}}^{v}$. This point is not smoothed by the deformation in Step I, and so also defines a toric fixed point in the toric surface $\widetilde{X}_{\mathscr{P}}^{v}$ that we continue to denote by $x_{ij}$. We denote by $(V_{ij},\partial V_{ij})$ the analytic germ of the surface pair $(\widetilde{X}_{\mathscr{P}}^{v},\partial\widetilde{X}_{\mathscr{P}}^{v})$ around $x_{ij}$.
Since the triangles $\Delta_{ij}$ have equal lattice height and lattice base, it follows from [11, §7.3] that $(V_{ij},\partial V_{ij})$ is the germ of the origin in an affine toric variety of the form
@@
with $n\geq 1$ and $\mathrm{gcd}(a,n)=1$. If $n=1$, then $(V_{ij},\partial V_{ij})$ is the germ of the origin in $(\mathbb{A}^{2},\{uv=0\})$. If $n>1$, then $V_{ij}$ is an elementary T-singularity as in [74, Proposition 3.10], also known as a Wahl singularity. Using the change of variables $x=u^{n}$, $y=v^{n}$, $z=uv$, this singularity can be equivalently described as $\raisebox{1.99997pt}{$(xy=z^{n})$}\left/\raisebox{-1.99997pt}{$\frac{1}{n}(1,-1,a)$}\right$.}, and admits $\mathbb{Q}$-Gorenstein smoothings by [74, Definition 3.7]. More precisely, its universal $\mathbb{Q}$-Gorenstein deformation space is irreducible, one-dimensional, and given by
$\raisebox{1.99997pt}{$(xy=z^{n}+t)$}\left/\raisebox{-1.99997pt}{$\frac{1}{n}(1,-1,a,0)$}\right.$}\,.
@@
By taking $\{z=0\}$ as deformation of $\partial V_{ij}$, we view this smoothing of $V_{ij}$ as a one-parameter deformation of the pair $(V_{ij},\partial{V}_{ij})$. If $n=1$, we consider the one-parameter deformation of $(\mathbb{A}^{2}_{u,v},\{uv=0\})$ given by $(\mathbb{A}^{2}_{u,v},\{uv=t\})$, that is by just smoothing the normal crossing toric divisor. Thus, for every $n$, we obtain a natural one-parameter $\mathbb{Q}$-Gorenstein deformation of $(V_{ij},\partial V_{ij})$, whose general fiber is the germ of a smooth point with a smooth connected divisor.